# Tensor Attention Training: Provably Efficient Learning of Higher-order Transformers

### v1

@@ F.6 Gradient computation in almost linear time by low rank tensor approximation
  • Let $k_{3}=n^{o(1)}$. We know that $\mathsf{F}_{a}(x)$ has approximate low rank representation $U_{3},V_{3},W_{3}\in\mathbb{R}^{k_{3}}$, let $\widetilde{\mathsf{F}}_{a}(x)$ denote $U_{3}(V_{3}\oslash W_{3})^{\top}$.
  • Let $k_{4}=n^{o(1)}$. We know that $\mathsf{F}_{b}(x)$ has approximate low rank representation $U_{4},V_{4},W_{4}\in\mathbb{R}^{k_{4}}$, let $\widetilde{\mathsf{F}}_{b}(x)$ denote $U_{4}(V_{4}\oslash W_{4})^{\top}$.
@@
$\frac{\mathsf{Loss}(x)}{\mathrm{d}x}=\operatorname{vec}(A_{1}^{\top}\mathsf{F}(x)(A_{2}\otimes A_{3}))$
@@
$\operatorname{vec}(\underbrace{A_{1}^{\top}}_{{d\times n}}\underbrace{\widetilde{\mathsf{F}}(x)}_{{n\times n^{2}}}\underbrace{(A_{2}\otimes A_{3})}_{{n^{2}\times d}})=$ $~{}\operatorname{vec}(\underbrace{A_{1}^{\top}}_{{d\times n}}\underbrace{\widetilde{\mathsf{F}}(x)}_{{n\times n^{2}}}\underbrace{(A_{2}^{\top}\otimes A_{3}^{\top})^{\top}}_{{n^{2}\times d}})$
@@
where the first step is due to Fact C.4, the second step is because of Claim C.20 and Fact C.12, and the last step follows Fact C.13
The above computation takes $n^{1+o(1)}d+O(d^{3})$ time. So, overall time complexity is still $n^{1+o(1)}$.
@@ G.2 Main result for lower bound
Suppose that $\mathsf{I}_{d}\in\mathbb{R}^{d\times d\times d}$ is an identity tensor. Also suppose that the input matrices to Lemma 6.2 are $Q,K_{1},K_{2},V_{1},V_{2}$. And we set $A_{1}=Q$, $A_{2}=K_{1}$,$A_{3}=K_{2}$, $A_{4}=V_{1}$,$A_{5}=V_{2}$, $Y=I$, and $X=\lambda\cdot\underbrace{\mathrm{mat}(\mathsf{I}_{d})}_{d\times d^{2}}$, with some $\lambda\in[0,1]$. Let $f:[0,1]\to\mathbb{R}$ be defined in Lemma G.1 where $H$ is the matrix $A_{1}(A_{2}\oslash A_{3})^{\top}$, so that $M_{\lambda}$ is the matrix $\exp(A_{1}X(A_{2}\otimes A_{3})^{\top})$ by Fact C.8. It follows from Lemma G.1 and $d=\Theta(\log n)$ that

### v2

@@ F.6 Gradient computation in almost linear time by low rank tensor approximation
  • Let $k_{3}=n^{o(1)}$. We know that $\mathsf{F}_{a}(x)$ has approximate low rank representation $U_{3},V_{3},W_{3}\in\mathbb{R}^{n\times k_{3}}$, let $\widetilde{\mathsf{F}}_{a}(x)$ denote $U_{3}(V_{3}\oslash W_{3})^{\top}$.
  • Let $k_{4}=n^{o(1)}$. We know that $\mathsf{F}_{b}(x)$ has approximate low rank representation $U_{4},V_{4},W_{4}\in\mathbb{R}^{n\times k_{4}}$, let $\widetilde{\mathsf{F}}_{b}(x)$ denote $U_{4}(V_{4}\oslash W_{4})^{\top}$.
@@
$\frac{\mathrm{d}\mathsf{Loss}(x)}{\mathrm{d}x}=\operatorname{vec}(A_{1}^{\top}\mathsf{F}(x)(A_{2}\otimes A_{3}))$
@@
$\operatorname{vec}(\underbrace{A_{1}^{\top}}_{{d\times n}}\underbrace{\widetilde{\mathsf{F}}(x)}_{{n\times n^{2}}}\underbrace{(A_{2}\otimes A_{3})}_{{n^{2}\times d^{2}}})=$ $~{}\operatorname{vec}(\underbrace{A_{1}^{\top}}_{{d\times n}}\underbrace{\widetilde{\mathsf{F}}(x)}_{{n\times n^{2}}}\underbrace{(A_{2}^{\top}\otimes A_{3}^{\top})^{\top}}_{{n^{2}\times d^{2}}})$
@@
where the first step is due to Fact C.4, the second step is because of Claim C.20 and Fact C.11, and the last step follows Fact C.12.
The above computation takes $n^{1+o(1)}d+d^{3}n^{o(1)}$ time. So, overall time complexity is still $n^{1+o(1)}$.
@@ G.2 Main result for lower bound
Suppose that $\mathsf{I}_{d}\in\mathbb{R}^{d\times d\times d}$ is an identity tensor. Also suppose that the input matrices to Lemma 6.2 are $Q,K_{1},K_{2},V_{1},V_{2}$. And we set $A_{1}=Q$, $A_{2}=K_{1}$,$A_{3}=K_{2}$, $A_{4}=V_{1}$,$A_{5}=V_{2}$, $Y=I$, and $X=\lambda\cdot\underbrace{\mathrm{mat}(\mathsf{I}_{d})}_{d\times d^{2}}$, with some $\lambda\in[0,1]$. Let $f:[0,1]\to\mathbb{R}$ be defined in Lemma G.1 where $H$ is the matrix $A_{1}(A_{2}\oslash A_{3})^{\top}$, so that $M_{\lambda}$ is the matrix $\exp(A_{1}X(A_{2}\otimes A_{3})^{\top})$ by Fact C.19. It follows from Lemma G.1 and $d=\Theta(\log n)$ that